A truck moves from rest with a uniform acceleration of 2m/s^2 for the first 10sec, it then accelerates at a uniform rate of 1m/s^2 for another 15sec, it continues at a constant speed for 70sec an finally come to rest in 20sec by uniform retardation.
1
Expert's answer
2021-08-17T10:29:17-0400

Given:

"a_1=2\:\rm m/s^2"

"a_2=1\:\rm m/s^2"

"t_1=10\:\rm s"

"t_2=15\:\rm s"

"t_3=70\:\rm s"

"t_4=20\:\rm s"


The total distance

"d=d_1+d_2+d_3+d_4"

1.

"d_1=\frac{a_1t_1^2}{2}=\frac{2*10^2}{2}=100\:\rm m"

2.

"v_2=a_1t_1=2*10=20\:\rm m/s"

"d_2=v_2t_2+\frac{a_2t_2^2}{2}=20*15+\frac{1*15^2}{2}=412.5\:\rm m"

3.

"v_3=v_2+a_2t_2=20+1*15=35\:\rm m/s"

"d_3=v_3t_3=35*70=2450\:\rm m"

4.

"d_4=\frac{v_3+0}{2}t_4=\frac{35+0}{2}*20=350\:\rm m"

Finally,

"d=100+412.5+35+2450+350=3312.5\:\rm m"